# Agree to Disagree? A Meta-Evaluation of LLM Misgendering

### v1

@@ Abstract
Numerous methods have been proposed to measure LLM misgendering, including probability-based evaluations (e.g., automatically with templatic sentences) and generation-based evaluations (e.g., with automatic heuristics or human validation). However, it has gone unexamined whether these evaluation methods have convergent validity, that is, whether their results align. Therefore, we conduct a systematic meta-evaluation of these methods across three existing datasets for LLM misgendering. We propose a method to transform each dataset to enable parallel probability- and generation-based evaluation. Then, by automatically evaluating a suite of 6 models from 3 families, we find that these methods can disagree with each other at the instance, dataset, and model levels, conflicting on 20.2% of evaluation instances. Finally, with a human evaluation of 2400 LLM generations, we show that misgendering behaviour is complex and goes far beyond pronouns, which automatic evaluations are not currently designed to capture, suggesting essential disagreement with human evaluations. Based on our findings, we provide recommendations for future evaluations of LLM misgendering. Our results are also more widely relevant, as they call into question broader methodological conventions in LLM evaluation, which often assume that different evaluation methods agree. Our code and data will be made available at: https://github.com/ArjunSubramonian/meta-eval-llm-misgendering.
@@ 1 Introduction
While most of these studies agree in their goals and the pronoun sets they consider, they use different methods to quantify misgendering. Some studies examine misgendering in LLM generations, while others evaluate whether LLMs assign a higher probability to a sequence that shows correct pronoun use, from a controlled set of minimally different templatic sequences. While generation is typically harder to evaluate, both automatically (Novikova et al., 2017; Colombo et al., 2023) and with humans (Howcroft et al., 2020), probability-based evaluations (e.g., template-based evaluations) have also been criticized for being brittle (Seshadri et al., 2022; Selvam et al., 2023) and uncorrelated with downstream biases (Goldfarb-Tarrant et al., 2021). Nevertheless, they remain widely used (Goldfarb-Tarrant et al., 2023).
@@
A question that has gone unexamined thus far is whether results from generation-based and probability-based evaluations correspond with or diverge from each other, i.e., whether they have convergent validity (Subramonian et al., 2023). Convergent validity captures whether “the measurements obtained from a measurement model” (e.g., the results of a probability-based evaluation of LLM misgendering) “correlate with other measurements of the same construct” (e.g., the results of a generation-based evaluation of misgendering) that have been “obtained from measurement models for which construct validity has already been established” (Jacobs & Wallach, 2021). This is particularly important given that LLMs can be used in different ways, sometimes for ranking existing sequences (Chiu & Chen, 2021; Salazar et al., 2020), and sometimes for generation, as in the popular chat models of today (OpenAI, ; Anthropic, ; Perplexity, ). However, NLP papers evaluating misgendering are not always explicit about which setups they are trying to evaluate, and sometimes probability-based evaluations are motivated by a desire to evaluate misgendering during generation. In such cases, if probability-based evaluations do not accurately estimate the propensity of models to misgender in real-world generations (e.g., due to the artificial nature of templates, limited prediction choices for the [MASK] token), they lack ecological validity, as they are “artificial [situations which do not] properly [reflect] broader real-world phenomenon” (Olteanu et al., 2019). In this paper, we thus set out to comprehensively and systematically compare evaluations of misgendering across four pronouns (he, she, they, xe).
For our meta-evaluation, we first transform three existing datasets to measure misgendering into parallel versions for probability- and generation-based evaluation (§4). By automatically evaluating a suite of 6 models from 3 families, we find that these methods disagree with each other on 20.2% of evaluation instances, and 24.2% of instances for the neopronoun xe (§5). This suggests that probability- and generation-based evaluations lack convergent and ecological validity depending on the application context. Next, with a human evaluation of 2400 LLM generations (§6), we show that misgendering behaviour goes beyond pronouns (McNamarah, 2021), which automatic generation-based evaluations do not currently capture (e.g., avoidance of pronoun usage, meta-discourse about pronouns, extraneous gendered terms). This suggests that automatic generation-based evaluations can inherently disagree with human evaluations. Finally, we come up with recommendations for future evaluations of LLM misgendering, such as critically considering the deployment context and recognizing the contextual nature of the appropriateness of gendered terms (§7). Our results are pertinent to other subfields of NLP (e.g., evaluations of stereotypes, linguistic acceptability), where probability-based LLM evaluations aim to measure factors that matter in generation contexts.
@@ 2 Related Work
Measuring LLM misgendering. A few works have contributed evaluations for LLM misgendering. Dev et al. (2021) present author-crafted templates to measure correct pronoun prediction for subjects with different names and pronouns, while Hossain et al. (2023) build on this with a more extensive set of templates, explicit pronoun declarations, and diverse pronoun cases for [MASK]. Gautam et al. (2024a) also use probability-based evaluation, but use implicit pronoun declarations, up to two subjects, and no personal names. In contrast, Ovalle et al. (2023) propose an automatic generation-based evaluation for misgendering in single-subject contexts. In this paper, we conduct a meta-evaluation of the agreement of the above-mentioned probability- and generation-based evaluations of LLM misgendering, and like Ovalle et al. (2023), include human evaluation in addition. We provide an overview of each dataset with the task and example inputs in Figure 1.
@@
Meta-evaluations of LM bias. Various probability-based evaluations (e.g., masked token, pseudo-log-likelihood) and generated text-based evaluations (e.g., distribution, classifier, lexicon) have been proposed for measuring bias (Gallegos et al., 2024). In response, some prior research has explored the lack of agreement between different LM bias evaluation methods. For example, several works have highlighted the inconsistency of probability-based bias measurements using templates (Delobelle et al., 2022; Seshadri et al., 2022; Selvam et al., 2023), and the unreliability of intrinsic bias metrics to measure application bias (Goldfarb-Tarrant et al., 2021; Cao et al., 2022). Moreover, templatic sentences may be poorly conceptualized (Blodgett et al., 2021) and can lack diversity, and comparisons of the probability of contrasting sentences do not capture the actual likelihood of models generating the sentences (Gallegos et al., 2024). In this paper, we focus on the agreement of likelihood, lexicon, and human-based measurements of misgendering with existing datasets.
@@
Lum et al. (2024) study disagreements between templatic “trick tests” (i.e., acontextual probability-based evaluations designed to elicit model bias) and realistic LLM use cases. They find that templatic “trick tests” are not predictive of bias in long-form text evaluations (i.e., story generation, user personas, ESL learning exercises). Similar to their work, we contribute a more tightly-coupled method to transform “trick test” datasets in a way that enables parallel probability- and generation-based evaluation of misgendering. We frame our work from a measurement modelling perspective, broadening conceptualizations of misgendering beyond pronouns, and examining convergent and ecological validity. Similarly, Goldfarb-Tarrant et al. (2023) discuss how the operationalization of bias measurements can be disconnected from how practitioners conceptualize bias, and Harvey et al. (2024) examine poor evaluation validity when metrics are disconnected from deployment contexts.
@@
Meta-evaluations of LMs in other contexts. Prior work has explored the extent to which direct LLM probabilities correlate with metalinguistic judgments (Hu & Levy, 2023; Song et al., 2025), and have found that metalinguistic judgments are not consistent indicators of model capabilities (Hu & Levy, 2023). Elangovan et al. (2025) examine how human uncertainty can affect measurements of the agreement of human and automatic evaluations.
@@ 3.3 Generation-Based Evaluation
In generation-based evaluation, the model receives a context sequence about a subject that surfaces a base pronoun of the subject (see Figure 2). The model then generates a completion sequence for the context. We say that the model misgenders the subject if it uses a pronoun to refer to the subject such that . For automatic evaluation of misgendering in these generations, we use the heuristic from Ovalle et al. (2023), i.e., choosing to be the first pronoun in the completion. Such heuristic functions are prone to error since pronoun generations can be about a different referent. Therefore, in Section 6, we validate this heuristic by manually annotating generations for misgendering. We provide further relevant details about and notation for the two evaluation formats in Appendix C.
In summary, probability-based evaluations assess whether LLMs assign a higher probability to templatic sequences that show correct pronoun use, from a controlled set of minimally different sequences with alternative pronouns. In contrast, generation-based evaluations measure the extent to which LLMs demonstrate correct pronoun use in open-ended generations. While one would expect LLMs to be more likely to generate sequences that they assign a higher probability, there can be deviations in the results of probability- and generation-based evaluations (e.g., due to LLMs being unlikely to generate templatic sequences, the autoregressive nature of decoding).
@@ 4 Experimental Setup
Below, we describe the models and datasets we use for our meta-evaluation. Since these datasets were originally designed for only one type of evaluation format (either generation-based or probability-based evaluation), we transform each dataset to support the other format. This creates a tightly-coupled, fairer comparison of the two methods, so that we can better understand inconsistencies between them. Additional details are provided in Appendix D.
@@ 4.3 Converting Generation-Based to Probability-Based Evaluations
To convert a generation-based evaluation dataset into a probability-based evaluation dataset , we transform each context and generation pair into a template , as shown in Figure 2. We first truncate such that there is only a single pronoun, and replace it with a [MASK] token, to create . Then, we concatenate to form . In Appendix D.4, we outline practical challenges we encountered with conversion.
5 Agreement between Probability-Based and Generation-Based Evaluations
@@ 5.2 Results
RUFF. Similar to Gen-MISGENDERED, Gen-RUFF also displays instance-level variation across generations, and disagreement in results between probability- and generation-based evaluation. In contrast to Gen-MISGENDERED, the methods tend to disagree the most with respect to when the subject uses they, with the largest disparity for Llama-8B. However, with respect to and , the methods have a low but higher agreement for they compared to other pronouns. This could be attributed to RUFF templates not containing personal names, which seem to be more polarizing in their gendered associations for LLMs.
@@ 6 Human Evaluation
Methodology. Two authors, both experts in English-language pronoun usage, annotated a total of 2400 generations – 25 pre-[MASK] and 25 post-[MASK] generations per ground-truth pronoun, for all models and the two datasets. Each generation was annotated for whether: (1) the ground-truth pronoun is correctly used, (2) misgendering occurs, or (3) no pronoun is generated. See Appendix G for the full annotation schema, and some observations made during annotation. In addition, we noted when models introduced extraneous gendered words such as “man,” “girl,” etc. On a sample of 200 instances that both authors annotated, agreement was 96% for pronoun labelling, and 98% for extraneous gendered information.
@@ 8 Conclusion and Future Work
In this work, we comprehensively compare generation-based and probability-based evaluations of misgendering in LLMs, by adapting three existing misgendering datasets for a parallel meta-evaluation. Our results show that these two evaluation approaches do not always converge, with disagreements on roughly 20% of instances. Through human evaluation, we show that misgendering is multifaceted and goes beyond just incorrect pronoun use. To address this, we recommend using community-grounded, holistic definitions of misgendering. More broadly, our empirical findings highlight the need for deliberate, reliable, and ecologically valid evaluation protocols. These findings are relevant beyond misgendering, in other subfields of NLP where relying on probability-based evaluations alone may fail to capture phenomena that occur in open-ended generation, and vice versa. We discuss the limitations of our work in Appendix A.
@@
We document our experimental details (e.g., hardware setup, runtime, decoding hyperparameters) in Appendix D, and we report our results over five generations for each dataset instance. Our code and data will be made available.
Acknowledgments and Disclosure of Funding

### v2

@@ Abstract
Numerous methods have been proposed to measure LLM misgendering, including probability-based evaluations (e.g., automatically with templatic sentences) and generation-based evaluations (e.g., with automatic heuristics or human validation). However, it has gone unexamined whether these evaluation methods have convergent validity, that is, whether their results align. Therefore, we conduct a systematic meta-evaluation of these methods across three existing datasets for LLM misgendering. We propose a method to transform each dataset to enable parallel probability- and generation-based evaluation. Then, by automatically evaluating a suite of 6 models from 3 families, we find that these methods can disagree with each other at the instance, dataset, and model levels, conflicting on 20.2% of evaluation instances. Finally, with a human evaluation of 2400 LLM generations, we show that misgendering behaviour is complex and goes far beyond pronouns, which automatic evaluations are not currently designed to capture, suggesting essential disagreement with human evaluations. Based on our findings, we provide recommendations for future evaluations of LLM misgendering. Our results are also more widely relevant, as they call into question broader methodological conventions in LLM evaluation, which often assume that different evaluation methods agree. Our code and data are available at: https://github.com/ArjunSubramonian/meta-eval-llm-misgendering.
@@ 1 Introduction
While these studies mostly agree in their goals and the pronoun sets they consider, they use different methods to quantify misgendering. Some studies examine misgendering in LLM generations, while others evaluate whether LLMs assign a higher probability to a sequence that shows correct pronoun use, from a controlled set of minimally different templatic sequences. While generation is typically harder to evaluate, both automatically (Novikova et al., 2017; Colombo et al., 2023) and with humans (Howcroft et al., 2020), probability-based templatic evaluations have also been criticized for being brittle (Seshadri et al., 2022; Selvam et al., 2023) and uncorrelated with downstream biases (Goldfarb-Tarrant et al., 2021). Nevertheless, they remain widely used (Goldfarb-Tarrant et al., 2023).
@@
A question that has gone unexamined thus far is whether results from generation-based and probability-based evaluations correspond with or diverge from each other, i.e., whether they have convergent validity (Jacobs & Wallach, 2021; Subramonian et al., 2023). This is particularly important given that LLMs can be used in different ways, sometimes for ranking existing sequences (Salazar et al., 2020; Chiu & Chen, 2021), and sometimes for generation, as in the popular chat models of today (OpenAI, ; Anthropic, ; Perplexity, ). However, NLP papers evaluating misgendering are not always explicit about which setups they are trying to evaluate (Goldfarb-Tarrant et al., 2023), and sometimes probability-based evaluations are motivated by a desire to evaluate misgendering during generation (Hossain et al., 2023; Gautam et al., 2024a). In such cases, if probability-based evaluations do not accurately estimate the propensity of models to misgender in real-world generations (e.g., due to the artificial nature of templates, limited prediction choices), they lack ecological validity, as they are “artificial [situations which do not] properly [reflect] broader real-world phenomenon” (Olteanu et al., 2019). In this paper, we thus set out to comprehensively and systematically compare evaluations of misgendering.
For our meta-evaluation, we first transform three existing datasets to measure misgendering into parallel versions for probability- and generation-based evaluation across four pronouns (he, she, they, xe) (§4). By automatically evaluating a suite of 6 models from 3 families, we find that these methods disagree with each other on 20.2% of evaluation instances, and 24.2% of instances for the neopronoun xe (§5). This suggests that probability- and generation-based evaluations lack convergent and ecological validity depending on the application context. Next, with a human evaluation of 2400 LLM generations (§6), we show that misgendering behaviour goes beyond pronouns (McNamarah, 2021), which automatic generation-based evaluations do not currently capture (e.g., pronoun avoidance, pronoun meta-discourse, extraneous gendered terms). This suggests that automatic generation-based evaluations can inherently disagree with human evaluations. Finally, we come up with recommendations for future evaluations of LLM misgendering, including critically considering the deployment context and recognizing the contextual nature of appropriate gendered terms (§7).
@@ 2 Related Work
Measuring LLM misgendering. A few works have contributed evaluations for LLM misgendering. Dev et al. (2021) present author-crafted templates to measure correct pronoun prediction for subjects with different names and pronouns, while Hossain et al. (2023) build on this with a more extensive set of templates, explicit pronoun declarations, and diverse pronoun cases for [MASK]. Gautam et al. (2024a) also use probability-based evaluation, but use implicit pronoun declarations, up to two subjects, and no personal names. In contrast, Ovalle et al. (2023) propose an automatic generation-based evaluation for misgendering in single-subject contexts. In this paper, we conduct a meta-evaluation of the agreement of the above-mentioned probability- and generation-based evaluations of LLM misgendering (see Figure 1 for an overview), and we include human evaluation as well.
@@
Meta-evaluations of LLM bias. Various probability-based evaluations (e.g., masked token, pseudo-log-likelihood) and generated text-based evaluations (e.g., distribution, classifier, lexicon) have been proposed for measuring bias (Gallegos et al., 2024). In response, some prior research has explored the lack of agreement between different LM bias evaluation methods. For example, several works have highlighted the inconsistency of probability-based bias measurements using templates (Delobelle et al., 2022; Seshadri et al., 2022; Selvam et al., 2023), and the unreliability of intrinsic bias metrics to measure application bias (Goldfarb-Tarrant et al., 2021; Cao et al., 2022). Moreover, templatic sentences may be poorly conceptualized and lack diversity (Blodgett et al., 2021), and comparisons of the probability of contrasting sentences do not capture the actual likelihood of models generating the sentences (Gallegos et al., 2024). In this paper, we focus on the agreement of likelihood, lexicon, and human-based measurements of misgendering with existing datasets.
@@
Lum et al. (2024) study disagreements between templatic “trick tests” (i.e., acontextual probability-based evaluations designed to elicit model bias) and realistic LLM use cases, finding that templatic “trick tests” are not predictive of bias in long-form text evaluations (i.e., story generation, user personas, ESL learning exercises). Parallel to their work, we contribute a more tightly-coupled method to transform datasets in a way that enables parallel probability- and generation-based evaluation of misgendering. In contrast to trick tests, however, the probability-based evaluations we consider measure more contextual, extrinsic bias. Similar to the measurement modelling perspective we take in our work, Goldfarb-Tarrant et al. (2023) discuss how the operationalization of bias measurements can be disconnected from how practitioners conceptualize bias, and Harvey et al. (2024) examine poor evaluation validity when metrics are disconnected from deployment contexts.
@@
Meta-evaluations of LLMs in other contexts. Prior work has explored the extent to which direct LLM probabilities correlate with metalinguistic judgments (Hu & Levy, 2023; Song et al., 2025), and have found that metalinguistic judgments are not consistent indicators of model capabilities (Hu & Levy, 2023). Elangovan et al. (2025) examine how human uncertainty can affect measurements of the agreement of human and automatic evaluations.
@@ 3.3 Generation-Based Evaluation
In generation-based evaluation, the model receives a context sequence about a subject that surfaces a pronoun (see Figure 2). The model then generates a completion sequence for the context. We say that the model misgenders the subject if it uses a pronoun to refer to the subject such that . For automatic evaluation of misgendering in these generations, we use the heuristic from Ovalle et al. (2023), i.e., choosing to be the first pronoun in the completion. Such heuristic functions are prone to error since pronoun generations can be about a different referent. Therefore, in Section 6, we validate this heuristic by manually annotating generations for misgendering. We provide further relevant details about and notation for the two evaluation formats in Appendix C.
In summary, probability-based evaluations assess whether LLMs assign a higher probability to templatic sequences that show correct pronoun use, from a controlled set of minimally different sequences with alternative pronouns. In contrast, generation-based evaluations measure the extent to which LLMs demonstrate correct pronoun use in open-ended generations. While one would expect LLMs to be more likely to generate sequences that they assign a higher probability, there can be deviations in the results of probability- and generation-based evaluations, e.g., due to LLMs being unlikely to generate templatic sequences and the autoregressive nature of decoding.
@@ 4 Experimental Setup
Below, we describe the models and datasets we use for our meta-evaluation. Since these datasets were originally designed for only one type of evaluation format (either generation- or probability-based evaluation), we transform each dataset to support the other format. This creates a tightly-coupled, fairer comparison of the two methods, so that we can better understand inconsistencies between them. Additional details are provided in Appendix D.
@@ 4.3 Converting Generation-Based to Probability-Based Evaluations
To convert a generation-based evaluation dataset into a probability-based evaluation dataset , we transform each context and generation pair into a template , as in Figure 2. We first truncate such that there is only one pronoun, and replace it with a [MASK] token, to create . Then, we concatenate to form . In Appendix D.4, we outline practical challenges we encountered with conversion.
5 Agreement between Probability- and Generation-Based Evaluations
@@ 5.2 Results
RUFF. Similar to Gen-MISGENDERED, Gen-RUFF also displays instance-level variation across generations, and disagreement in results between probability- and generation-based evaluation. In contrast to Gen-MISGENDERED, the methods tend to disagree the most with respect to when the subject uses they, with the largest disparity for Llama-8B. However, with respect to and , the methods have a low but higher agreement for they compared to other pronouns. The discrepancies between RUFF and MISGENDERED could be attributed to RUFF templates not containing personal names, which seem to be more polarizing in their gendered associations for LLMs.
@@ 6 Human Evaluation
Methodology. Two authors, both experts in English-language pronoun usage and their sociolinguistic norms, annotated a total of 2400 generations – 25 pre-[MASK] and 25 post-[MASK] generations per ground-truth pronoun, for all models and the two datasets. Annotations were conducted without reference to model outputs, which ensured that human evaluations of misgendering were performed solely based on contextual understanding. Each generation was annotated for whether: (1) the ground-truth pronoun is correctly used, (2) misgendering occurs, or (3) no pronoun is generated. See Appendix G for the full annotation schema, and some observations made during annotation. In addition, we noted when models introduced extraneous gendered words such as “man,” “girl,” etc. On a sample of 200 instances that both authors annotated, raw agreement was 96% for pronoun labelling, and 98% for extraneous gendered information.
@@ 8 Conclusion and Future Work
In this work, we comprehensively compare generation-based and probability-based evaluations of misgendering in LLMs, by adapting three existing misgendering datasets for a parallel meta-evaluation. Our results show that these two evaluation approaches do not always converge, with disagreements on roughly 20% of instances. Through human evaluation, we show that misgendering is multifaceted and goes beyond just incorrect pronoun use. To address this, we recommend using community-grounded, holistic definitions of misgendering. More broadly, our empirical findings highlight the need for deliberate, reliable, and ecologically valid evaluation protocols. These findings are relevant beyond misgendering, in other subfields of NLP (e.g., evaluations of stereotypes, linguistic acceptability) where relying on probability-based evaluations alone may fail to capture phenomena that occur in open-ended generation, and vice versa. We discuss limitations of our work in Appendix A.
@@
We document our experimental details (e.g., hardware setup, runtime, decoding hyperparameters) in Appendix D, and we report our results over five generations for each dataset instance. Our code and data are available at https://github.com/ArjunSubramonian/meta-eval-llm-misgendering.
Acknowledgements